OpenStudy (anonymous):

6(4y-3)=4(5y-4)-5 SHOW YOUR WORK!!

OpenStudy (anonymous):

first the distibute the 6 into the parentheses on the left sided.Then on the right sides dsitribute the 4 into the parentheses. It should become clear from there

OpenStudy (anonymous):

this problem is not clear to me at all

OpenStudy (anonymous):

you asked questions similar to this one already, and they have been explained clearly, go back to them and you will see that this problem is simlar

OpenStudy (anonymous):

6(4y-3)=4(5y-4)-5 Next distribute: 24y-18=20y-16-5 You can solve from here

OpenStudy (anonymous):

i got -6/2